OpenStudy (anonymous):

A train needs 20 seconds to go through a 100 yard tunnel. The very same train going at a constant speed needs 50 seconds to go through a 330 yard tunnel. What is the length of the train?

ganeshie8 (ganeshie8):

you need to setup two equations... which part you need help with ?

ganeshie8 (ganeshie8):

call, length of train = x speed of trian = s

OpenStudy (lgbasallote):

actually...it's doable with one equation only

ganeshie8 (ganeshie8):

use the relation, distance = speed*time

OpenStudy (lgbasallote):

since speed is constant

ganeshie8 (ganeshie8):

hmm

OpenStudy (lgbasallote):

something like distance/time = distance/time

OpenStudy (maheshmeghwal9):

Is the speed of train in both cases is equal? I mean it is given r nt?

OpenStudy (maheshmeghwal9):

r=or*

OpenStudy (maheshmeghwal9):

this question needs a clarification.

ganeshie8 (ganeshie8):

100+x = 20*s --------(1) 330+x = 50*s--------(2)

OpenStudy (lgbasallote):

something like this \[\huge \frac{100 + L}{20} = \frac{330 + L}{50}\]

OpenStudy (lgbasallote):

where L is length of train

OpenStudy (lgbasallote):

you guys do the rest ;)

ganeshie8 (ganeshie8):

oh..

OpenStudy (maheshmeghwal9):

But my doubt is other @lgbasallote :/

ganeshie8 (ganeshie8):

proportions are cute :D

OpenStudy (maheshmeghwal9):

I think we can't take same speed in both cases as it is nt stated.

OpenStudy (lgbasallote):

speed is constant...that means speed is the same at both instances

OpenStudy (maheshmeghwal9):

speed is constant in second case only but wt about the first case. question needs a clarification.

OpenStudy (lgbasallote):

*sigh*

OpenStudy (lgbasallote):

im not willing to do a grammar tutorial right now

OpenStudy (maheshmeghwal9):

but it is a type of physics question too physics do nt recommend this @lgbasallote ur solution is right 100% but question need clarification @CosmicStorm :/